OpenStudy (anonymous):

1. (4x^2+12x-16/2x+10) / (6x+24/x^2+9x+20) 2. (1/x-3) +(4/x) / (4/x) - (1/x-3) 3. (1/36)- (1/x^2) / (1/6)(1/x)

OpenStudy (anonymous):

Please Help me. I keep getting them wrong. Thank you <3

OpenStudy (anonymous):

\[\frac{\frac{4x^{2}+12x-16}{2x+10}}{\frac{6x+24}{x^{2}+9x+20}}\]\[=\frac{4(x^{2}+3x-4)}{2(x+5)}[\frac{x^{2}+9x+20}{6(x+4)}]\]can u simplify now?

OpenStudy (anonymous):

Yes i can Thank You...

OpenStudy (anonymous):

\[\frac{\frac{1}{x-3}+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x-3}}\]multiply numerator and denominator with x(x-3).

OpenStudy (anonymous):

Oh okay .

OpenStudy (anonymous):

as for the 3rd question, its quite similiar with ur 1st 1. you need to flip the denominator into multiplying the inverse, and then find the common denominator.

OpenStudy (anonymous):

Okay. I'm going to try again.

OpenStudy (anonymous):

Thank You.